OpenStudy (anonymous):

5. The quotient of a number and 6 is -108. Write an equation to solve for the number.

OpenStudy (anonymous):

you have to write the sentence out as a formula. the quotient means to divide, "a number" is just what you call x, and the word "is" means equals. so then we write it all out x/6=-108

OpenStudy (anonymous):

\[\frac{ x }{ 6 }=-108\]

OpenStudy (anonymous):

now you have to get the x by itself, so you have to multiply both sides by 6. and you get x=-108(6) which is x=-648

OpenStudy (anonymous):

Thank you !